OpenStudy (anonymous):

2 McDonald's Quarter Pounders and 3 burger king whoppers with cheese contain 250 milligrams of cholesterol. 3 quarter pounders and one whooper with cheese contain 353 milligrams of cholesterol . What is the cholesterol for each item?

OpenStudy (campbell_st):

Q = Quarter pounder W = Whopper set up 2 equations 2Q + 3W = 250 (1) 3Q + W = 353 (2) rewrite equation 2 W = 353 - 3Q substitute into equation (1) 2Q + 3(353 - 3Q) = 250 2Q - 9Q + 1059 = 250 -7Q = - 809 \[Q = \frac{809}{7}\] when you have Q substitute into equation 1 to find W and the answer is a bit strange.... with a very high Q and small W